Goodness of Fit Test 5. Aresearcher wanted to determine bicycle deaths were uniformly distributed over the days of the week. She randomly selected 200 deaths by bicycles, recorded the days they happened with the following data Monday Fuesday Wednesday Thursday Friday Saturday 16 35 16 28 34 30 s there evidence to believe that bicycle deaths occur with equal frequency with respect to day of the week at a 0.05 level of significance
Answer #1

5 )

The hypothesis are :

H0 : Bicycle deaths occur with equal frequency with respect to day.

H1: Bicycle deaths occur with unequal probability with respect to day.

The test statistic is

(O - E)2

O = Observed frequency

E = Expected frequency

E = N* p

All days have same probability. So p = 1/7 and N= 200.

Hence expected frequency = E = 200 * 1/7 = 28.57 for each day.

O E (O- E)
16 28.57 5.530
35 28.57 1.447
16 28.57 5.530
28 28.57 0.011
34 28.57 1.032
41 28.57 5.408
30 28.57 0.072
Total 19.031

So, test statistic is

chi ^{2} = 19.031

df = Number of days - 1 = 7 -1 = 6 , alpha = 0.05

In the chi square probability table , we look across df = 6 and probability 0.05 .

So the critical value = chi ^{2}_{alpha ,df}=chi ^{2}_{0.05,6}=12.592

Here calculated value of chi ^{2} > critical value of chi ^{2}

Hence we reject null hypothesis.

Conclusion : There is no evidence to believe that bicycle deaths occur with unequal frequency with respect to day.